Slab 21 (first layer in the full stack) was moved, after the shower program to the TB24/2 area.

We tested magnetic fields of 0, 1 and 0.5 T (in this order). we also took some reference data without beam.

The chip/spill/dif configuration was the same than previously.

We didn't see any clear failure/loss of performance during visual inspection on the web cam & online monitor, after a total of ~20 hours of data taking.

All power supplies were placed into the hut to avoid magnetic fields affecting them. This required very long cables and larger values of the input voltages to compensate the drop on currents.

- DIF: 4,11V au lieu de 3,3V - ASU: 5,7V au lieu de 5V - GDCC: 7,19V au lieu de 5,5V - CCC: 4.11 au lieu de 3,3V, 5,4V au lieu de 5V et -5,2V au lieu de -5V

-- AdrianIrles - 2017-10-10

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2017-10-10 - AdrianIrles
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    CALICE All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback